OpenStudy (anonymous):

Using Galilean transformations describe the motion of a falling stone in a moving train relative to a stationary observer on the ground and draw a sketch showing the motion

OpenStudy (roadjester):

Galileo transformation: \(x'=x-vt\) \(y'=y\) \(z'=z\) \(t'=t\)

OpenStudy (roadjester):

\[\large v\space'=dx'/dt'={d(x\space')\over d(t\space')}={d(x-vt)\over d(t)}={dx-vdt\over dt}\]